Suppose that the riskless interest rate is zero. Suppose that a stock is priced at $100, and one year from now will be at eitI want to know how to get the answers. Thanks
Answer #1

Expected price 1 year from now=Sum(probability*stock price)=0.80*130+0.20*70=118

Expected return=ln(price 1 year from now/price now)=ln(118/100)=16.55%

ln is natural logarithm or logarithm to the base e